OpenStudy (anonymous):

derived Is it Correct? please check this answer \[f(x) = \sqrt[3]{1+\cos(2x)/1-\cos(2x)}\] \[f'(x)= \frac{ \sqrt[3]{1+\sin2/1-\sin2}\ }{ 3* \sqrt[3]{(1+\cos2x/1-\cos2x)^{2}} }\]

OpenStudy (anonymous):

i think not . you have to use both power and quotient rules. your question is this ? \[\Large f(x)= \sqrt[3]{\frac{(1+\cos(2x)}{1-\cos(2x)}}\]

OpenStudy (experimentx):

I would ask wolfram. and use this http://www.codecogs.com/latex/eqneditor.php for latex

OpenStudy (anonymous):

is very large @experimentX

OpenStudy (experimentx):

yep ... perfect for practice to develop your skills!!

OpenStudy (anonymous):

ok thanks i will try

OpenStudy (anonymous):